# Optimizing Temperature Distributions for Training Neural Quantum States using Parallel Tempering

### v2

@@ I Introduction
The flexibility of artificial neural networks (ANNs) has made them an attractive choice for parameterized wavefunctions that can be trained using variational Monte Carlo (VMC) methods, such as stochastic reconfiguration (SR) [14, 15] and minimum-step stochastic reconfiguration [16], to approximate ground states [17]. An ANN used to represent a quantum state is now commonly referred to as a Neural Quantum State (NQS). State-of-the-art results can be achieved by carefully choosing the ANN architecture to reproduce the symmetries of the Hamiltonion of interest[18, 19]. For recent reviews of NQSs and their applications, see Ref. [20, 21, 22]. Despite the success and promise of NQSs, training these networks to approximate the ground state can still be resource intensive [23], with the primary culprits being frustration and the sign structure of the ground state [24, 25, 26].
@@
The manuscript is organized as follows. In Sec. II, we describe our methodology, including the ANN architectures, variational algorithms, and PT method used. In Sec. III, we present our results for the two classes of problems we study, the Precipice problem and the $J_{1}$-$J_{2}$ model. The Precipice problem has an energy landscape that drives the training to a local minimum from which it is very difficult to escape and serves as a useful toy model to demonstrate how our temperature optimization can improve performance. Because of its simplicitly, we study this model with only a Restricted Boltzmann Machine (RBM) [32, 33]. For the $J_{1}$-$J_{2}$ model, we use both a RBM and a feedforward [34] network architecture in order to illustrate the usefulness of our methods on different architectures. In Sec. IV, we discuss our results and conclusions.
@@ II.1 Wavefunction Representation
where the set $\left\{|x\rangle\right\}_{x}$ is the computational basis for $n$ qubits [^1] We pick the convention that for the Pauli $z$ operator on a single qubit, the computational basis satisfies $\sigma^{z}|0\rangle=|0\rangle$ and $\sigma^{z}|1\rangle=-|1\rangle$.. The function $\psi_{x}(\alpha)$ is the output of an ANN with parameters $\left\{\alpha_{k}\right\}_{k}$ given the input $x$. In this work, we focus on two types of ANN’s, the Restricted Boltzmann Machine (RBM) and a three layer feedforward neural network. For a RBM with $n$ visible nodes and $m$ hidden nodes, the parameterized function is given by:
@@ III.1 Precipice problem
In contrast, in Fig. 1(b) we show how the swap probability for all pairs of replicas converges to a non-zero value. The exact non-zero value depends on the number of replicas used: a higher value is reached as more replicas are used. Unlike the case without temperature optimization, the algorithm is now able to find the ground state. For the particular case illustrated in Fig. 1(b), this coincides with a ‘blip’ in the swap probability between the 0 and 1 replicas, where the configuration at $T_{\min}$ finds a lower energy and its configuration is swapped with the configuration at $T=0$. This can be seen in Fig. 2 where the expectation value of the Hamiltonian for the zero temperature replica is shown to finally reach the ground state at that time.
In Fig. 2, we show how the frequency of finding the ground state depends on the number of replicas. For 6 replicas and less, the swap probabilities are too low to allow the algorithm to find the ground state with the fixed number of updates we performed. As the number of replicas is increased, the number of updates needed to find the ground state is reduced, but the benefit of having more replicas diminishes once the swap probability becomes too large. This is consistent with what is observed in more ‘traditional’ studies using PT where an optimal swap probability is observed [43].
@@ III.2 $J_{1}$-$J_{2}$ Model with RBMs
For our first case, we focus on a $4\times 5$ system with $J_{1}=1$ and $J_{2}=0.695$, training with an RBM. This is close to the point where the gap between the first excited state and ground state is smallest, near $J_{2}=0.7$, and where we find the greatest difficulty in training a RBM.
@@
We use a RBM ansatz with $n=20$ visible units and $m=40$ hidden units for our variational wavefunction. This ansatz is likely not expressive enough to fully capture the ground state; we have not observed energies reaching the ground state in any of our simulations. However, it is expressive enough to capture states with energies below the first excited state, which we term as success throughout the rest of this section.
@@
In Fig. 3 we measure the likelihood of reaching an energy below the first excited state with a single replica and with PT with 6 replicas. In the case of PT, we consider the two cases of allowing temperature optimization throughout the training and holding the temperature fixed throughout the training. The $J_{1}$-$J_{2}$ problem is inherently difficult for a single replica RBM system to solve without any PT, with a success rate of only 3% $\pm$ 3% after $2.5\times 10^{4}$ updates. PT without any temperature optimization improves the overall success rate to 30% $\pm$ 9%, demonstrating a clear improvement. Once we allow for temperature optimization, we achieve a significantly higher success rate of 65% $\pm$ 9%. In Fig. 3, we show how the temperature $1/\beta$ changes over time for a typical run where we allow the temperatures to optimize. We find that our initial polynomial distribution of temperatures is relatively good and only small adjustments are made throughout the run. This demonstrates that optimizing temperatures can increases success probability even for systems that are initially well tuned.
@@
The reason for this improvement can be seen by examining the swap probabilities between each of the different replicas for a typical run, as shown in Fig. 4. When the temperatures are allowed to optimize, the swap probabilities converge near 0.3, although the values are still noisy due to the limited sample size and number of updates. For transferring into the zero replica (which contains the pure target Hamiltonian), the probability of swapping is a binary 0 or 1, based on whether replica 1 has a lower target energy. When we take the running average in Fig. 4, this corresponds to the averaged probability moving up for update regions where there is more transfer into the zero replica. This move toward high probabilities for swapping into/out of the zero replica is particularly pronounced at $10^{4}$ updates and $1.6\times 10^{4}$ updates. The latter of these update states corresponds precisely with when the energy of the zero replica begins to lower, eventually dipping below the first excited state. In contrast, when the temperatures are not allowed to optimize as in Fig. 4, the probabilities of swapping never converge toward any value. Furthermore, the probability of swapping between the 2nd and 3rd replicas remains at essentially 0 throughout the entirety of the simulation. This leads to an inability for any configurations from replicas 3 to 5 to ever make their way to the 0 replica, which likely leads to the simulation never reaching energies below the first excited state. While these results are for two specific runs of PT for this problem, they are representative for how the algorithm achieves success for a majority of our simulations. This shows that optimizing temperatures allows the PT algorithm to perform better by ensuring that states are mixed in from the higher temperatures in the distribution at a steady rate.
We further find that the temperature optimization algorithm is relatively robust to the initial choice of temperature distribution. In Fig. 5, we explore a linear initial distribution of temperatures, $T_{k}=T_{\min}+(T_{\max}-T_{\min})\left(\frac{k-1}{n_{\mathrm{R}}-2}\right)$, where $k=1,\dots,n_{\mathrm{R}}-1$. The probability of success at $J_{2}=0.695$ nearly exactly matches what we see for the polynomial distribution in Fig. 3. Without optimizing the temperatures, we see a success rate of $0.22\pm 0.082$. Once we allow the temperatures to optimize, the success rate improves to $0.68\pm 0.093$. In Fig. 5, the $\beta$ values in a representative example run adjust significantly away from their initial linear distribution within just approximately $10^{3}$ updates. This distribution of temperatures closely matches the distribution that we see when starting with an initially polynomial distribution. From these results we conclude that our algorithm is robust to choices of the initial temperature distributions.
@@
We do however observe a relatively strong dependence on the choice of minimum and maximum temperatures, which are held fixed throughout the simulation. In Fig. 6, we vary the maximum temperature $T_{\max}$ for a system with $n_{R}=6$ and an initial polynomial distribution of temperatures and note how the success rate changes. We see that for maximum temperatures greater $\geq 10$, the success rate for simulations where the temperature distribution is optimized holds steady at $\approx 0.60$. For simulations without temperature optimization, the success rate varies from $\approx 0.20$ to $0.40$. Since setting the maximum temperature also affects the initial temperature distribution, this variation is expected.
However as the maximum temperature decreases below 10, the success rate for the simulations drastically increases, with a choice of $T_{\max}=2$ giving a success rate of $0.88$ for both simulations with and without temperature optimization. This shows that the choice of $T_{\max}=20$ explored in the earlier figures was not optimal for maximizing success. Therefore the success frequency of our algorithm is still highly sensitive to the initial choice of the maximum temperature. Nevertheless, the temperature optimization does provide robust improvement to the success rate even for suboptimal choices of initial temperature, which can speed up the process of finding viable solutions. Future work will focus on automating the selection of the $T_{\max}$ parameter.
@@ III.3 $J_{1}$-$J_{2}$ Model with Feedforward Networks
As a final test, we study the $J_{1}$-$J_{2}$ on a square lattice with periodic boundary conditions for a $6\times 6$ system. In this case, we choose $J_{1}=1$ and $J_{2}=0.62$. This is close to the point where the energy gap between the first excited state and the ground state is smallest, near $J_{2}=0.58$. For this study, we use a three-layer feedforward neural network instead of a RBM, and we denote the number of hidden nodes in each layer by $(h_{1},h_{2},h_{3})$. This network is more expressive than the RBMs used in Sec. III.2 and is intended to test our method on a more expressive architecture. For this system size, exact sampling is computationally prohibitive and prevents us from evaluating the energy of the NQS exactly, so we employ a different criterion to define success. We take a running mean of the energy and standard deviation over 50 steps and if the average energy reported is beneath $-70.07405$ (computed using DMRG [44, 45, 46, 47] with a bond dimension of 2048), by more than a third of the mean standard deviation, we freeze the parameters and sample at that step for $10^{3}$ steps with 1024 samples each. Each of these sampling runs gives a sample mean energy and a sample standard deviation, from which we can calculate a mean of the means energy and a standard error of the mean. If the mean of the mean energies is one standard error below the first excited state energy, then we flag the simulation as a possible success. When the simulation terminates we again sample, this time for $10^{3}$ steps with 512 samples each in order to confirm the success of the simulation. We find that all simulations that are flagged successful at intermediate points in the simulation are confirmed successful with this final test.
@@
We show the success of PT using the $(48,24,24)$ feedforward network in Fig. 7. We observe a clear advantage to using PT over not using PT in terms of success frequency, and we also begin to see a separation between the simulations with and without temperature optimization after approximately $10^{4}$ update steps. After $1.5\times 10^{4}$ steps, the increase in success rate for the case without temperature optimization slows down and possibly saturates near $60\%$, suggesting that the system can get stuck in local minima. This is in contrast to the case with temperature optimization where a properly tuned temperature distribution can help it escape from such local minima and succeed with a rate near $90\%$.
@@ IV Conclusions
In this work, we demonstrated the impact of the replica temperature distribution on the effectiveness of PT to help train an ANN to represent a low-energy quantum state. Training with PT already shows increased success rates over training with a single replica, and by adaptively adjusting the temperature distribution such that all pairs of replicas swap with equal probability we reach even higher success frequencies at quicker rates with negligible computational cost. We attribute this improvement to better escape rates from local minima. These results are more pronounced for problems where the training is difficult. This can arise from deep local minima in the energy potential, which was illustrated by the Precipice problem, but also when the number of parameters gets larger for more expressive networks, as was demonstrated by our different network sizes for the $6\times 6$ $J_{1}$-$J_{2}$ model. We expect that the difficulties encountered for the $J_{1}$-$J_{2}$ problem will only be exacerbated at larger problem sizes, so we can expect temperature optimization to be crucial in that case.
We have focused on neural networks that are generic to demonstrate the effectiveness of our proposed methodology. Replacing the RBM or feedforward networks with more state-of-the-art networks adopted for NQS such as GCNN’s [18, 19] is straightforward. We expect the advantages of PT to arise when the training is stymied by local minima from which the variational algorithm is slow to escape, and the PT replica swaps provide a mechanism for escape, for which optimizing the temperature distribution is important to be effective.

### v3

@@ I Introduction
The flexibility of artificial neural networks (ANNs) has made them an attractive choice for parameterized wavefunctions that can be trained using variational Monte Carlo (VMC) methods, such as stochastic reconfiguration (SR) [14, 15] and minimum-step stochastic reconfiguration [16], to approximate ground states [17]. An ANN used to represent a quantum state is now commonly referred to as a Neural Quantum State (NQS). State-of-the-art results can be achieved by carefully choosing the ANN architecture to reproduce the symmetries of the Hamiltonian of interest[18, 19]. For recent reviews of NQSs and their applications, see Ref. [20, 21, 22]. Despite the success and promise of NQSs, training these networks to approximate the ground state can still be resource intensive [23], with the primary culprits being frustration and the sign structure of the ground state [24, 25, 26].
@@
We emphasize that training with PT does not improve the expressivity of the NQS used, so the lowest energy that could in principle be achieved is not changed. Success of the approach is instead measured in terms of how long it takes the training to reach a desired level of accuracy. For this reason, our work focuses on the number of updates it takes to reach a target energy for the same NQS, and we do not use state-of-the-art NQS’s that can achieve the lowest known energies. In practice, choosing the NQS that most efficiently approximates the ground state and the training algorithm that most effectively reaches that approximation should go hand in hand.
The manuscript is organized as follows. In Sec. II, we describe our methodology, including the ANN architectures, variational algorithms, and PT method used. In Sec. III, we present our results for the two classes of problems we study, the Precipice problem and the $J_{1}$-$J_{2}$ model. The Precipice problem has an energy landscape that drives the training to a local minimum from which it is very difficult to escape and serves as a useful toy model to demonstrate how our temperature optimization can improve performance. Because of its simplicitly, we study this model with only a Restricted Boltzmann Machine (RBM) [32, 33]. For the $J_{1}$-$J_{2}$ model, we use both an RBM and a feedforward [34] network architecture in order to illustrate the usefulness of our methods on different architectures. In Sec. IV, we discuss our results and conclusions.
@@ II.1 Wavefunction Representation
where the set $\left\{|x\rangle\right\}_{x}$ is the computational basis for $n$ qubits [^1] We pick the convention that for the Pauli $z$ operator on a single qubit, the computational basis satisfies $\sigma^{z}|0\rangle=|0\rangle$ and $\sigma^{z}|1\rangle=-|1\rangle$.. The function $\psi_{x}(\alpha)$ is the output of an ANN with parameters $\left\{\alpha_{k}\right\}_{k}$ given the input $x$. In this work, we focus on two types of ANN’s, the Restricted Boltzmann Machine (RBM) and a three layer feedforward neural network. For an RBM with $n$ visible nodes and $m$ hidden nodes, the parameterized function is given by:
@@ III.1 Precipice problem
In contrast, in Fig. 1(b) we show how the swap probability for all pairs of replicas converges to a nonzero value. The exact nonzero value depends on the number of replicas used: a higher value is reached as more replicas are used. Unlike the case without temperature optimization, the algorithm is now able to find the ground state. For the particular case illustrated in Fig. 1(b), this coincides with a ‘blip’ in the swap probability between the 0 and 1 replicas, where the configuration at $T_{\min}$ finds a lower energy and its configuration is swapped with the configuration at $T=0$. This can be seen in Fig. 2 where the expectation value of the Hamiltonian for the zero temperature replica is shown to finally reach the ground state at that time.
In Fig. 2, we show how the frequency of finding the ground state depends on the number of replicas. For 6 replicas and fewer, the swap probabilities are too low to allow the algorithm to find the ground state with the fixed number of updates we performed. As the number of replicas is increased, the number of updates needed to find the ground state is reduced, but the benefit of having more replicas diminishes once the swap probability becomes too large. This is consistent with what is observed in more ‘traditional’ studies using PT where an optimal swap probability is observed [43].
@@ III.2 $J_{1}$-$J_{2}$ Model with RBMs
For our first case, we focus on a $4\times 5$ system with $J_{1}=1$ and $J_{2}=0.695$, training with an RBM. This is close to the point where the gap between the first excited state and ground state is smallest, near $J_{2}=0.7$, and where we find the greatest difficulty in training an RBM.
@@
We use an RBM ansatz with $n=20$ visible units and $m=40$ hidden units for our variational wavefunction. This ansatz is likely not expressive enough to fully capture the ground state; we have not observed energies reaching the ground state in any of our simulations. However, it is expressive enough to capture states with energies below the first excited state, which we term as success throughout the rest of this section.
@@
In order to quantify the advantage of our method, we use the likelihood of the simulation reporting an energy beneath the first excited state as opposed to the mean energy achieved. This is because we are not trying to gauge the expressibility of the wavefunction but the computational efficiency of PT. In Fig. 3 we measure this likelihood and compare the case of a single replica to the case of PT with 6 replicas. In the case of PT, we consider the two cases of allowing temperature optimization throughout the training and holding the temperature fixed throughout the training. The $J_{1}$-$J_{2}$ problem is inherently difficult for a single replica RBM system to solve without any PT, with a success rate of only 3% $\pm$ 3% after $2.5\times 10^{4}$ updates. PT without any temperature optimization improves the overall success rate to 30% $\pm$ 9%, demonstrating a clear improvement. Once we allow for temperature optimization, we achieve a significantly higher success rate of 65% $\pm$ 9%. In Fig. 3, we show how the temperature $1/\beta$ changes over time for a typical run where we allow the temperatures to optimize. We find that our initial polynomial distribution of temperatures is relatively good and only small adjustments are made throughout the run. This demonstrates that optimizing temperatures can increase success probability even for systems that are initially well tuned.
@@
The reason for this improvement can be seen by examining the swap probabilities between each of the different replicas for a typical run, as shown in Fig. 4. When the temperatures are allowed to optimize, the swap probabilities converge near 0.3, although the values are still noisy due to the limited sample size and number of updates. For transferring into the zero replica (which contains the pure target Hamiltonian), the probability of swapping is a binary 0 or 1, based on whether replica 1 has a lower target energy. When we take the running average in Fig. 4, this corresponds to the averaged probability moving up for update regions where there is more transfer into the zero replica. This move toward high probabilities for swapping into/out of the zero replica is particularly pronounced at $10^{4}$ updates and $1.6\times 10^{4}$ updates. The latter of these update states corresponds precisely with when the energy of the zero replica begins to lower, eventually dipping below the first excited state. In contrast, when the temperatures are not allowed to optimize as in Fig. 4, the probabilities of swapping never converge toward any value. Furthermore, the probability of swapping between the 2nd and 3rd replicas remains at essentially 0 throughout the entirety of the simulation. This leads to an inability for any configurations from replicas 3 to 5 to ever make their way to the 0 replica, which likely leads to the simulation never reaching energies below the first excited state. While these results are for two specific runs of PT for this problem, they are representative of how the algorithm achieves success for a majority of our simulations. This shows that optimizing temperatures allows the PT algorithm to perform better by ensuring that states are mixed in from the higher temperatures in the distribution at a steady rate.
We further find that the temperature optimization algorithm is relatively robust to the initial choice of temperature distribution. In Fig. 5, we explore a linear initial distribution of temperatures, $T_{k}=T_{\min}+(T_{\max}-T_{\min})\left(\frac{k-1}{n_{\mathrm{R}}-2}\right)$, where $k=1,\dots,n_{\mathrm{R}}-1$. The probability of success at $J_{2}=0.695$ nearly exactly matches what we see for the polynomial distribution in Fig. 3. Without optimizing the temperatures, we see a success rate of $0.22\pm 0.082$. Once we allow the temperatures to optimize, the success rate improves to $0.68\pm 0.093$. In Fig. 5, the $\beta$ values in a representative example run adjust significantly away from their initial linear distribution within just approximately $10^{3}$ updates. This distribution of temperatures closely matches the distribution that we see when starting with an initially polynomial distribution. From these results, we conclude that our algorithm is robust to choices of the initial temperature distributions.
@@
We do, however, observe a relatively strong dependence on the choice of minimum and maximum temperatures, which are held fixed throughout the simulation. In Fig. 6, we vary the maximum temperature $T_{\max}$ for a system with $n_{R}=6$ and an initial polynomial distribution of temperatures and note how the success rate changes. We see that for maximum temperatures greater $\geq 10$, the success rate for simulations where the temperature distribution is optimized holds steady at $\approx 0.60$. For simulations without temperature optimization, the success rate varies from $\approx 0.20$ to $0.40$. Since setting the maximum temperature also affects the initial temperature distribution, this variation is expected.
However, as the maximum temperature decreases below 10, the success rate for the simulations drastically increases, with a choice of $T_{\max}=2$ giving a success rate of $0.88$ for both simulations with and without temperature optimization. This shows that the choice of $T_{\max}=20$ explored in the earlier figures was not optimal for maximizing success. Therefore, the success frequency of our algorithm is still highly sensitive to the initial choice of the maximum temperature. Nevertheless, the temperature optimization does provide robust improvement to the success rate even for suboptimal choices of initial temperature, which can speed up the process of finding viable solutions. Future work will focus on automating the selection of the $T_{\max}$ parameter.
@@ III.3 $J_{1}$-$J_{2}$ Model with Feedforward Networks
As a final test, we study the $J_{1}$-$J_{2}$ on a square lattice with periodic boundary conditions for a $6\times 6$ system. In this case, we choose $J_{1}=1$ and $J_{2}=0.62$. This is close to the point where the energy gap between the first excited state and the ground state is smallest, near $J_{2}=0.58$. For this study, we use a three-layer feedforward neural network instead of an RBM, and we denote the number of hidden nodes in each layer by $(h_{1},h_{2},h_{3})$. This network is more expressive than the RBMs used in Sec. III.2 and is intended to test our method on a more expressive architecture. For this system size, exact sampling is computationally prohibitive and prevents us from evaluating the energy of the NQS exactly, so we employ a different criterion to define success. We take a running mean of the energy and standard deviation over 50 steps and if the average energy reported is beneath $-70.07405$ (computed using DMRG [44, 45, 46, 47] with a bond dimension of 2048), by more than a third of the mean standard deviation, we freeze the parameters and sample at that step for $10^{3}$ steps with 1024 samples each. Each of these sampling runs gives a sample mean energy and a sample standard deviation, from which we can calculate a mean of the means energy and a standard error of the mean. If the mean of the mean energies is one standard error below the first excited state energy, then we flag the simulation as a possible success. When the simulation terminates we again sample, this time for $10^{3}$ steps with 512 samples each in order to confirm the success of the simulation. We find that all simulations that are flagged successful at intermediate points in the simulation are confirmed successful with this final test.
@@
We show the success of PT using the $(48,24,24)$ feedforward network in Fig. 7. We include the mean energies of these simulations in Appendix D, but we again find success probability to be a more informative metric. We observe a clear advantage to using PT over not using PT in terms of success frequency, and we also begin to see a separation between the simulations with and without temperature optimization after approximately $10^{4}$ update steps. After $1.5\times 10^{4}$ steps, the increase in success rate for the case without temperature optimization slows down and possibly saturates near $60\%$, suggesting that the system can get stuck in local minima. This is in contrast to the case with temperature optimization where a properly tuned temperature distribution can help it escape from such local minima and succeed with a rate near $90\%$.
@@ IV Conclusions
In this work, we demonstrated the impact of the replica temperature distribution on the effectiveness of PT to help train an ANN to represent a low-energy quantum state. Training with PT already shows increased success rates over training with a single replica, and by adaptively adjusting the temperature distribution such that all pairs of replicas swap with equal probability, we reach even higher success frequencies at quicker rates with negligible computational cost. We attribute this improvement to better escape rates from local minima. These results are more pronounced for problems where the training is difficult. This can arise from deep local minima in the energy potential, which was illustrated by the Precipice problem, but also when the number of parameters gets larger for more expressive networks, as was demonstrated by our different network sizes for the $6\times 6$ $J_{1}$-$J_{2}$ model. We expect that the difficulties encountered for the $J_{1}$-$J_{2}$ problem will only be exacerbated at larger problem sizes, so we can expect temperature optimization to be crucial in that case.
We have focused on neural networks that are generic to demonstrate the effectiveness of our proposed methodology. Replacing the RBM or feedforward networks with more state-of-the-art networks adopted for NQS such as GCNN’s [18, 19] is straightforward. While it remains to be tested, we expect PT to be useful with larger architectures due to the difficulty inherent in their training. The FermiNet architecture [48], for example, typically utilizes nearly $\mathcal{O}(10^{6})$ parameters and $\mathcal{O}(10^{5})$ training iterations alongside a pretraining procedure. PT may enable either the number of parameters or total training iterations to be decreased for larger networks like this. While larger networks make finding local minima close to the global minimum of the training landscape easier, they also make finding the global minimum exponentially more difficult [49]. If all of the local minima are a good approximation of the ground state, then increasing the network size alone would be a successful strategy. However, it is also possible that the global minimum is a qualitatively different state from any of the local minima, in which case finding any local minimum may not be sufficient. We therefore believe PT to be particularly useful in cases where finding the global minimum with higher efficiency is needed. When PT is used in tandem with large networks, we expect it to increase the probability of converging to the global minimum or to a minimum close to the limit of the network’s expressibility. It may also decrease the number of training iterations required to converge and allow us to use smaller networks to achieve similar energies. All three of these advantages in efficiency become more important with larger networks.